OpenStudy (anonymous):

Limit of (x^7-1)/(x^6+1) as x approaches infinity.

OpenStudy (bahrom7893):

infinity

OpenStudy (bahrom7893):

because 7>6

OpenStudy (anonymous):

SO when the Degree of the numerator is bigger than the degree of the denominater. it's infinity???

OpenStudy (anonymous):

If I follow that rule, isn't it also possible that the answer doesn't exist? how do you know which is which?

OpenStudy (anonymous):

\[\lim _{x \rightarrow \infty} \frac{x^7-1}{x^6+1}=\frac{\infty}{\infty}\]

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

you have an inderterminate form so use l'hopitals \[f'(x)=7x^6\] \[g'(x)=6x^5\] still infinity, take f'' , g'' \[f''(x)=42x^5\] \[g''(x)=30x^4\] eventually the bottom will disappear leaving you with a x on top which will get you inifinity

OpenStudy (anonymous):

ok heres the work i got: lim x--> o-o(infinity) it is an "indeterminate form of o-o/o-o so use the largest power denominator and multiply the denominator and numerator by it so you'll get like: lim x --> o-o x^7-1 * 1/x^6 / x^6+1 * 1/x^6 -Now remembering the reciprocal power rule here (if an infinity is over a real number it goes to 0) you're going to get x/1 So what is lim x--> o-o of x/1?

OpenStudy (bahrom7893):

No there's an easier way to explain this..

OpenStudy (bahrom7893):

Basically since the top is higher power... it approaches infinity faster than the bottom

OpenStudy (anonymous):

ya, very true, but what if her teacher wants the work, and i think it helps to understand why

OpenStudy (anonymous):

you can also use l'hopitals rule. \[\lim_{x \rightarrow \infty} \frac{ f(x) }{ g(x) } = \lim_{x \rightarrow \infty} \frac{ f'(x) }{ g'(x) }\] where f(x) = x^7-1 and g(x) = x^6 +1 f'(x) = 7x^6 and g'(x) = 6x^5. \[\frac{ 7*x^6 }{ 6x^5 } = (7/6)*x = infinity\]

OpenStudy (bahrom7893):

ohh

OpenStudy (zarkon):

\[\lim _{x \rightarrow \infty} \frac{x^7-1}{x^6+1}=\lim _{x \rightarrow \infty} \frac{x^6}{x^6}\frac{x-1/x^6}{1+1/x^6}=...\]

OpenStudy (zarkon):

l'hospitals is overkill for such a simple problem

OpenStudy (anonymous):

eh, I think like a calculator, and a calculator would solve it wth l'hopitals

OpenStudy (anonymous):

because it's much easier to program if (both = infnity) , then take derivatives..

OpenStudy (anonymous):

Ohhh. Gotcha. Thank you. I like the I'hopital rule. I've never heard of it. Thank you.

OpenStudy (anonymous):

Ok guys. how about helping me with one more?

OpenStudy (zarkon):

you should be careful if you are not familiar with L'Hospitals rule. It is easy to get wrong answers if certain conditions are not met

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty } x ^{3}-5x ^{2}\]

OpenStudy (zarkon):

Notice... \[x ^{3}-5x ^{2}=x^2(x-5)\]

OpenStudy (anonymous):

yesss

OpenStudy (zarkon):

as \(x\to\infty\) you have \(x^2\to\infty\) and \(x-5\to\infty\)

OpenStudy (anonymous):

So the answer is infinity?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

and if it were aprouching negative infinity, would it be negative infinity?

OpenStudy (zarkon):

in this case ...yes

OpenStudy (anonymous):

Ok thanks! Eee, I just got helped by Data ^_^ lol

OpenStudy (zarkon):

you are welcome